the sum of a,b and c is 14. if 9a + 8b + 6c = 101, what is the value of 2a + b -c?
Answer by tommyt3rd(5050)

9a + 8b + 6c = 101 and a+b+c=14 so if we multiply every term by 7 we get: 7a+7b+7c=98
and if we subtract the result from the first equation we get 2a+b-c=3

so it's 3

:)